# Blameless Users in a Clean Room: Defining Copyright Protection for Generative Models

### v1

@@ 3.3.2 $k$-NAF does not prevent full reconstruction
In more detail, we give a family of models that are $k$-NAF with respect to pure noise, yet which enable a user to reconstruct the dataset verbatim. For any $\ell\geq 1$, let $\mathsf{coin}_{\ell}(\cdot|x)$ be uniform over $\{0,1\}^{\ell}$ for all prompts $x$. As a generative model, $\mathsf{coin}_{\ell}(\cdot|\cdot)$ is clearly a “safe” instantiation of $\mathsf{safe}_{c}$ for any copyrighted work $c$.
@@
  • There exists a user $\mathsf{u}$ such that for all $D$ and $k>0$: $\mathsf{u}$ makes $\operatorname*{poly}(L,1/k)$ black-box queries to $p_{k,D}$ and outputs $D$ with probability $>0.99$.
The proof is deferred to Appendix B.2.
@@ 4 The interaction between user and model
Let $\mathsf{u}$ be a (randomized) algorithm called the user, $p$ be a model, and $\mathsf{aux}\in\{0,1\}^{*}\cup\{\bot\}$ be an auxiliary input. We denote by $\mathsf{u}^{p}(\mathsf{aux})$ the algorithm $\mathsf{u}$ run with input $\mathsf{aux}$ and black-box access to $p$. The result is a work $z\in\mathcal{W}$ distributed as $z\leftarrow\mathsf{u}^{p}(\mathsf{aux})$. Together with a training algorithm $\mathsf{Train}$ and a dataset $D$, this process induces probability measure $\tau$ over $\mathcal{W}$.
@@
The probability is taken over the randomness of the algorithms $\mathsf{Train}$, $\mathsf{u}$, and $p$.
@@ 5 Tainted models enable copying
This section defines what it means for a generative model to be tainted. Tainted models let users copy training data of which they have no knowledge. Tainted models blatantly fail to prevent copying. If the goal is to prevent copying, being tainted is as bad as it gets. A tainted model—one produced by a tainted training algorithm (Definition 5.1)—doesn’t prevent copyright infringement. It enables it! Thus, we get our next desideratum. (We will see that NAF fails this test but Definition 7.6 passes.)
Desideratum 2.
We say a training algorithm $\mathsf{Train}$ is tainted if there is a fixed algorithm $\mathsf{u}$ that copies work from the training dataset using only trained model and the unprotected ideas of the work in question.
@@
The order of quantifiers is important. The user $\mathsf{u}$ is fixed once and for all, and $D$ varies arbitrarily. The user cannot possibly be at fault: it cannot “know” a work in all possible datasets $D$.
@@ 5.1 Examples
Example 5.2 below describes a black-box transformation from any training algorithm to a tainted-version of that training algorithm. This example is used to prove Theorem 3.5.
@@
Algorithm 1 describes a tainted algorithm $\mathsf{Train}_{\mathsf{kv}}$, constructed from any training algorithm $\mathsf{Train}_{0}$ in a black-box way. In words, $\mathsf{Train}_{\mathsf{kv}}(D)$ trains a model $q_{0}\leftarrow\mathsf{Train}_{0}(D)$, and also builds a key-value store $I$. $I$ maps ideas $\mathsf{id}$ to the set $D_{\mathsf{id}}=\{w\in D:\mathsf{ideas}(w)=\mathsf{id}\}$. On prompt $x$, the model $q_{\mathsf{kv}}$ returns a random element of $I[x]=D_{x}$ if it is non-empty. Otherwise, it returns a generation sampled from $q_{0}(\cdot|x)$. It is easy to see that $\mathsf{Train}_{\mathsf{kv}}$ is tainted. Consider the user $\mathsf{u}^{p}(\mathsf{aux})$ that returns a sample from $p(\cdot|\mathsf{aux})$. Fix $D$ and $w\in D$, and let $p\leftarrow\mathsf{Train}_{\mathsf{kv}}(D)$. By construction, the user always outputs an element of $D_{\mathsf{ideas}(w)}$. Therefore, $\tau(\mathsf{SubSim}(D);\mathsf{ideas}(w))\geq\tau(D_{\mathsf{ideas}(w)};\mathsf{ideas}(w))=1$.
Parameters: Training algorithm $\mathsf{Train}_{0}$
Input: Data $D$
Output: Model $q_{\mathsf{kv}}$
Let $q_{0}\leftarrow\mathsf{Train}_{0}(D)$;
Initialize an empty key-value store $I$, whose keys are ideas $\mathsf{id}$ and values are sets of works $W\subset\mathcal{W}$;
for $w\in D$ do
       // add w to I[ideas(w)];
       $W\leftarrow I[\mathsf{ideas}(w)]$;
       $W\leftarrow W\cup\{w\}$;
       $I[\mathsf{ideas}(w)]\leftarrow W$;
end for
Let $q_{\mathsf{kv}}$ the conditional generative model which on prompt $x$ does the following:;
$W\leftarrow I[x]$;
if $W\neq\emptyset$ then
      Return $y\mathrel{{\leftarrow}\vbox{\hbox{\tiny\$}}}W$.
else
      Return $y\sim q_{0}(\cdot|x)$.
end if
Result: $q_{\mathsf{kv}}$
As a negative example, the following claim states that any training algorithm that always returns a single fixed model is not tainted, under a mild assumption.
@@ 5.2 Tainted NAF models
Corollary 5.4.
For any $\mathsf{ideas}$, algorithm $\mathsf{CP}^{*}$ given by Thm. 3.4 is tainted.
@@
Immediate from the statement of Thm. 3.5
@@ 6 Blameless copy protection: a definitional framework
We can’t hope to guarantee that a generative model never reproduces copyrighted work, even works it was not trained on. A malicious user can always induce copying. (Formally, if the set of in-copyright works $\mathcal{C}$ is non-empty, then there exists $\mathsf{u}$ that always infringes: $\tau(\mathsf{SubSim}(\mathcal{C});\bot)=1$.) Instead, we wish protect blameless users—users who don’t themselves induce infringement—from unwitting copying.
@@
Fix $\kappa>0$ and a class $\mathfrak{B}$ of blameless users. We say $\mathsf{Train}$ is $(\kappa,\mathfrak{B})$-copy protective if for all $\mathcal{C}\subseteq\mathcal{W}$, $D\in\mathcal{W}^{*}$, $\mathsf{aux}$, and $\mathsf{u}\in\mathfrak{B}$: $\tau(\mathsf{SubSim}(\mathcal{C});~{}\mathsf{aux})\leq\kappa$.
@@ 7 Defining clean-room copy protection
This section formalizes what it means to have access to a copyrighted work (Section 7.1) and instantiates the blameless copy protection framework (Section 7.2)—drawing inspiration from clean-room design. In copyright law, a clean room is “a process of producing a product under conditions guaranteeing independent design and foreclosing the possibility of copying.” [^4] David S. Elkins, NEC v. Intel: A Guide to Using “Clean Room” Procedures as Evidence, 10 Computer L.J. 453 (1990). https://repository.law.uic.edu/jitpl/vol10/iss4/1 The idea comes from cases involving reverse engineering. Roughly, a team is given a description of design specs of the product (ideas) but not the product itself (expression), and tasked with producing a compatible product. There is no access, as long as the team itself was not spoiled by prior familiarity with the product or its design. As such, the outcome is constructively non-infringing. Even substantial similarities will be the product of independent creation.
@@ 7.1 Scrubbing a dataset removes access
We need a way to say whether a work $w^{\prime}$ depends on or derives from another work $w$ a copyright-relevant way. Such dependencies are directed, with later works stemming from earlier works. The copyright dependency graph encodes this relation. It is assumed to capture the legal concept of access in the following sense: For purposes of copyright law, a dataset $D$ only gives access to a work $w$ if there is some $w^{\prime}\in D$ that stems from $w$.
@@
Definition 7.2 (The sets $\mathcal{C}_{D}$ and $\mathcal{C}_{-D}$).
We define $\mathcal{C}_{D}=\{c\in\mathcal{C}:\exists w\in D,(c,w)\in E\}$ as the set copyrighted works from which any work in $D$ stems. We denote its complement $\mathcal{C}_{-D}=\mathcal{C}\setminus\mathcal{C}_{D}$.
@@ 7.2 Clean training algorithms: copy protection for blameless users in a clean room
The user’s clean-room distribution
@@
Clean-room blamelessness and copy protection
We postulate that blameless users can control their risk of copying when using a model trained in a clean room. Given $\beta>0$, a blameless user can guarantee that they copy a work $c$ that was “outside the clean room” with probability at most $\beta$ [^5] The exact constant doesn’t matter. It should be small enough to be very unlikely to occur by chance ($\ll 2^{-10}$). It should be large enough to reflect that a relatively small amount of original expression is needed for copyright protection ($\gg 2^{-1000}$). Here’s a very crude justification for $\beta\approx 2^{-50}$. Estimates of the entropy of English text vary. Shannon estimated it at 0.6 to 1.3 bits per character (A–Z and space), based on a study of how well his English-speaking test subjects were able to predict the next character in a passage [Sha51]. With those estimates, just 77 to 167 characters of text contain $\approx$100 bits of entropy. Here is Shakespeare’s Sonnet 130 with the 77th and 167th characters highlighted. My mistress’ eyes are nothing like the sun;
Coral is far more red than her lips’ red;
If snow be white, why then her breasts are dun;
If hairs be wires, black wires grow on her head.
Acknowledging that I know nothing of the matter, somewhere in that range seems reasonable for the minimum amount of original expression needed to afford copyright protection. Infringement only requires substantial similarity, not perfect reproduction. If substantial similarity requires half of the 100 bits of minimum original expression to be reproduced, we get to $2^{-50}$.
  This should hold even if the user is exposed to $c$ in some other than through the model. [^6] One can weaken this assumption (making more users blameless) by restricting $\mathsf{aux}$ in Definitions 7.5 and 7.6. For example, by allowing $\mathsf{aux}$ to include $\mathsf{ideas}(c)$ but not original expression of works in $\mathcal{C}_{-D}\cup\{c\}$. This would undermine Desideratum 1, yielding qualitatively weaker protection. DP would still suffice (Theorem 8.2) with this change. People are constantly exposed to copyrighted works, yet we somehow to manage to avoid copying every day. Using a model trained without access to $c$ shouldn’t change that.
@@
For $0\leq\beta\leq 1$, a user $\mathsf{u}$ is $\beta$-blameless in the clean room ($\beta$-blameless) with respect to $D$, $\mathcal{C}$, $\mathsf{Train}$ if for all $c\in\mathcal{C}$ and all $\mathsf{aux}$:
@@
Otherwise, $\mathsf{u}$ is $\beta$-blameworthy.777 Equivalently, $\mathfrak{B}^{\mathsf{clean}}_{\beta}(D,\mathcal{C},\mathsf{ideas}):=\{\mathsf{u}:\forall c\in\mathcal{C},\forall\mathsf{aux},\tau_{-c}\bigl{(}\mathsf{SubSim}\bigl{(}\mathcal{C}_{-D}\cup\{c\}\bigr{)};~{}\mathsf{id}\bigr{)}\leq\beta\}.$
@@
Clean-room copy protection is not tainted
By Desideratum 2, a good definition of provable copy protection should exclude tainted training algorithms. The following theorem shows that clean-room copy protection passes this test, under a reasonable assumption on $\mathsf{ideas}$ and $\mathsf{SubSim}$. Roughly, the assumption is that there exist at least $n$ distinct ideas each of which can be expressed in $m\gg n$ ways satisfying a strong dissimilarity proprty: that for any distinct $w$ and $w^{\prime}$, the set of works substantially similar to both $w$ and $w^{\prime}$ is empty.
@@ 8 Differential privacy’s protection against copying
Many works have suggested that differential privacy (DP) should offer some sort of protection against copying if the training data are deduplicated [BLM20, HLJ+23, VKB23, EKHLM24, CKOX24]. This section turns the suggestion into a theorem. For background on DP [DMNS06], see Appendix C.
At a high level, we show that DP training blameless users from copyright infringement (in the sense of Definition 7.6) if the training dataset is golden, a copyright-deduplication condition defined below. If so, the probability of copying is at most $\approx e^{\varepsilon}\beta N_{D}$, where $\beta$ is a user’s probability of copying in the clean room and $N_{D}=|\mathcal{C}_{D}|$ is the number of copyrighted works to which $D$ grants access. To guarantee the risk is at most $\kappa$, the user sets $\beta\approx{\kappa}/{e^{\varepsilon}N_{D}}$. (This is pessimistic: see Remark 8.4.)
A dataset $D$ is golden if at most one item $w\in D$ stems from any copyrighted work $c$ [^8] We adapt and formalize this idea from [VKB23], where it appears as an informal condition that suffices for their analysis: “It is important that when we omit a data point $x$ it does not share copyrighted content with many other data points that were included in the training set.” We also borrow the term “golden dataset”, though [VKB23] uses it to refer to something entirely different: A dataset that is “carefully scrutinized to ensure that all material in it is not copyrighted or [is] properly licensed.” That condition doesn’t suffice, as even licensed derivates could cause unlicensed copying by a generative model. The same is true for NAF, despite their suggestion that it would yield a safe model.   That item could be $c$ itself or a derivative work. No protected original expression appears in more than one element of a golden dataset. For example, it can contain one copy or parody of the Abbey Road album cover (not both), and many parodies of the out-of-copyright Mona Lisa. Remark 8.5 explains why golden datasets will typically remain golden as the set of in-copyright works evolves over time.
@@
Let $G=(\mathcal{W},E)$ be a copyright dependency graph, and $\mathcal{C}\subseteq\mathcal{W}$ be the set of in-copyright works. A dataset $D$ is golden (with respect to $G$, $\mathcal{C}$) if for all $c\in\mathcal{C}$ sthere is at most one $w\in D$ such that $(c,w)\in E$.
@@
Fix $\mathcal{C}\subseteq\mathcal{W}$, $D$ and user $\mathsf{u}$ that is $\beta$-blameless (Def. 7.5). We must show that for all $\mathsf{aux}$, $\tau(\mathsf{SubSim}(\mathcal{C});~{}\mathsf{aux})\leq\kappa$. To simplify notation, we omit $\mathsf{aux}$ for the rest of the proof. Because $D$ is golden, $|D\setminus\mathsf{scrub}(D,c)|\leq 1$. Hence, datasets $\mathsf{scrub}(D,c)$ and $D$ are either equal or neighboring for all $c\in\mathcal{C}$. Applying DP and post-processing, we have $\tau(E)\leq e^{\varepsilon}\cdot\tau_{-c}(E)+\delta$ for all events $E$ and all $c\in\mathcal{C}$.
$\tau(\mathsf{SubSim}(\mathcal{C}))$ $\leq\tau(\mathsf{SubSim}(\mathcal{C}_{-D}))+\sum_{c\in\mathcal{C}_{D}}\tau(\mathsf{SubSim}(c))$ (union bound)
$\leq\beta+N_{D}\delta+e^{\varepsilon}\cdot\sum_{c\in\mathcal{C}_{D}}\tau_{-c}(\mathsf{SubSim}(c))$ (DP; Prop. 8.3 below)
$\leq(e^{\varepsilon}N_{D}+1)\beta+N_{D}\delta\leq\kappa$ ($foo\displaystyle\text{($\beta$-blameless)\phantom{foo}}\qed$
@@
Let $\mathsf{u}$ be $\beta$-blameless w.r.t. $D$, $\mathcal{C}$, $\mathsf{Train}$. For all $c\in\mathcal{C}_{-D}$, $\mathsf{aux}$: $\tau\left(\mathcal{C}_{-D};~{}\mathsf{aux}\right)\leq\beta.$
@@
Remark 8.4 (Improving the union bound).
The factor of $N_{D}=|\mathcal{C}_{D}|$ from the union bound is unreasonably pessimistic. Ideas sometimes differ so greatly that the risk of substantial similarity is essentially 0: a photorealistic bird will never resemble Dr. Seuss’s Cat in the Hat. If the user is targeting a fixed set of ideas $\mathsf{id}$ (conditioning $\tau$ on $\mathsf{id}$), one might instead take the union bound over $\mathcal{C}_{D}^{\mathsf{id}}=\{c\in\mathcal{C}_{D}:\mathsf{id}\in\mathsf{ideas}(\mathsf{SubSim}(c))\}$. We expect that typically $N_{D}^{\mathsf{id}}:=|\mathcal{C}_{D}^{\mathsf{id}}|\ll|\mathcal{C}_{D}|=N_{D}$.
@@
9 Discussion: provable copyright protection in practice?
What if copying occurs?

### v2

@@ 3.3.2 $k$-NAF does not prevent full reconstruction
Our next theorem shows that $k$-NAF may allow training data to be reconstructed, even if $k$ is arbitrarily small and independent of $x$. This is because the $k$-NAF guarantee does not compose across a user’s many queries, and each query may leak up to $k$ bits of training data. In more detail, we give a family of models that are $k$-NAF with respect to pure noise, yet which enable a user to reconstruct the dataset verbatim. For any $\ell\geq 1$, let $\mathsf{coin}_{\ell}(\cdot|x)$ be uniform over $\{0,1\}^{\ell}$ for all prompts $x$. As a generative model, $\mathsf{coin}_{\ell}(\cdot|\cdot)$ is clearly a “safe” instantiation of $\mathsf{safe}_{c}$ for any copyrighted work $c$. The proof is deferred to Appendix B.2.
@@
  • There exists a user $\mathsf{u}$ such that for all $D$ and $k>0$: $\mathsf{u}$ makes $\operatorname*{poly}(L,1/k)$ black-box queries to $p_{k,D}$ and outputs $D$ with probability >0.99>0.99.
@@ 4 The interaction between user and model
Let $\mathsf{u}$ be a (randomized) algorithm called the user, $p$ be a model, and $\mathsf{aux}\in\{0,1\}^{*}\cup\{\bot\}$ be an auxiliary input. In cryptography, auxiliary inputs allow one to define security guarantees even when an algorithm has instance-specific side information. In our setting, the algorithm is the user, and the side information represents everything the user knows that is external to the model. Often, we will take $\mathsf{aux}$ to be the non-copyrightable ideas of an in-copyright work: $\mathsf{aux}=\mathsf{ideas}(c)$.
We denote by $\mathsf{u}^{p}(\mathsf{aux})$ the algorithm $\mathsf{u}$ run with input $\mathsf{aux}$ and black-box access to $p$. The result is a work $z\in\mathcal{W}$ distributed as $z\leftarrow\mathsf{u}^{p}(\mathsf{aux})$. Together with a training algorithm $\mathsf{Train}$ and a dataset $D$, this process induces probability measure $\tau$ over $\mathcal{W}$, defined next.
@@
The probability is taken over the randomness of the algorithms $\mathsf{Train}$, $\mathsf{u}$, and $p$. Note that $\tau$ depends on $\mathsf{Train}$, $\mathsf{u}$, and $D$. The notation elides these dependencies to reduce clutter.
@@ 5 Tainted models enable copying
This section defines what it means for a generative model to be tainted. Tainted models let users copy training data of which they have no knowledge. We say a training algorithm $\mathsf{Train}$ is tainted if there is a fixed algorithm $\mathsf{u}$ that copies work from the training dataset using only trained model and unprotected ideas. The order of quantifiers is important: the user $\mathsf{u}$ is fixed while the dataset $D$ is arbitrary. Such a user cannot possibly be at fault—it cannot know a work in all possible datasets $D$.
@@
If the goal is to prevent copying, being tainted is as bad as it gets. Thus, we get our next desideratum. We will see that NAF fails this test but Definition 7.6 passes.
Desideratum 2.
@@ 5.1 Examples
As an example, we show that any training algorithm can be made tainted (Example 5.2). This is used to prove Theorem 3.5.
@@
Algorithm 3 describes a tainted algorithm $\mathsf{Train}_{\mathsf{kv}}$, constructed from any training algorithm $\mathsf{Train}_{0}$ in a black-box way. In words, $\mathsf{Train}_{\mathsf{kv}}(D)$ trains a model $q_{0}\leftarrow\mathsf{Train}_{0}(D)$, and also builds a key-value store $I$ mapping ideas $\mathsf{id}$ to the set $D_{\mathsf{id}}=\{w\in D:\mathsf{ideas}(w)=\mathsf{id}\}$. On prompt $x$, the model $q_{\mathsf{kv}}$ returns a random element of $I[x]=D_{x}$ if it is non-empty. Otherwise, it returns a generation sampled from $q_{0}(\cdot|x)$. It is easy to see that $\mathsf{Train}_{\mathsf{kv}}$ is tainted. Consider the user $\mathsf{u}^{p}(\mathsf{aux})$ that returns a sample from $p(\cdot|\mathsf{aux})$. Fix $D$ and $w\in D$, and let $p\leftarrow\mathsf{Train}_{\mathsf{kv}}(D)$. By construction, the user always outputs an element of $D_{\mathsf{ideas}(w)}$. Therefore, $\tau(\mathsf{SubSim}(D);\mathsf{ideas}(w))\geq\tau(D_{\mathsf{ideas}(w)};\mathsf{ideas}(w))=1$.
As a negative example, the following claim states that any training algorithm that always returns a single fixed model is not tainted, under the mild assumption that it is possible for two works to express the same ideas so differently that no work is similar to both.
@@ 5.2 Tainted NAF models
Corollary 5.4 (of Thm. B.5).
For any $\mathsf{ideas}$, the NAF algorithm $\mathsf{CP}^{*}$ given by Thm. 3.4 is tainted.
@@
Immediate from the statement of Theorem 3.5
@@ 6 Blameless copy protection: a definitional framework
We can’t hope to guarantee that a generative model never reproduces copyrighted work, even works it was not trained on. Malicious users can always induce copying. (Formally, if the set of in-copyright works $\mathcal{C}$ is non-empty, there exists $\mathsf{u}$ that always infringes: $\tau(\mathsf{SubSim}(\mathcal{C});\bot)=1$.) Instead, we wish to protect blameless users—who don’t themselves induce infringement—from unwitting copying.
@@
Fix $\kappa>0$ and a class $\mathfrak{B}$ of blameless users. We say $\mathsf{Train}$ is $(\kappa,\mathfrak{B})$-copy protective if for all blameless $\mathsf{u}\in\mathfrak{B}$ and all $\mathsf{aux}$, $D\in\mathcal{W}^{*}$, $\mathcal{C}\subseteq\mathcal{W}$:
$\tau(\mathsf{SubSim}(\mathcal{C});~\mathsf{aux})<\kappa.$
We place no a priori restriction on the auxiliary information $\mathsf{aux}$, which represents everything the user knows that is external to the model. This could include copyrighted material. For example, Alice might be writing a children’s book and also own a copy of Green Eggs and Ham. As long as she’s careful, Alice isn’t doing anything wrong and should still enjoy some protection.
@@ 7 Defining clean-room copy protection
This section formalizes what it means to have access to a copyrighted work (Section 7.1) and instantiates the blameless copy protection framework (Section 7.2)—drawing inspiration from clean-room design. In copyright law, a clean room is “a process of producing a product under conditions guaranteeing independent design and foreclosing the possibility of copying” [elkins1990nec]. The clean room keeps contamination out, like keeping dust out of a semiconductor lab. The idea comes from cases involving reverse engineering. Roughly, a team is given a description of design specs of the product (ideas) but not the product itself (expression), and tasked with producing a compatible product. There is no access, as long as the team itself was not spoiled by prior familiarity with the product or its design. As such, the outcome is constructively non-infringing. Even substantial similarities will be the product of independent creation.
Clean-room design is the inspiration for a counterfactual clean-room distribution where a user interacts with a model trained without access to certain in-copyright works (Definition 7.4). We quantify the blamelessness of a user as the probability $\beta$ that the user—in the clean-room distribution—would have produced something that is substantially similar to the excluded works (Definition 7.5). Clean-room copy protection provides a corresponding bound $\kappa$ on the probability of copying any work in the real distribution (Define 7.6), where $\kappa$ depends on $\beta$.
@@ 7.1 Scrubbing a dataset removes access
We need a way to say whether a work $w^{\prime}$ is a derivative of another work $w$ a copyright-relevant way. Such dependencies are directed, with later works stemming from earlier works. The copyright dependency graph encodes this relation. It is assumed to capture the legal concept of access in the following sense: For purposes of copyright law, a dataset $D$ only gives access to a work $w$ if there is some $w^{\prime}\in D$ that stems from $w$.
@@
Definition 7.2 ($\mathcal{C}_{D}$ and $\mathcal{C}_{-D}$).
We define $\mathcal{C}_{D}=\{c\in\mathcal{C}:\exists w\in D,(c,w)\in E\}$ as the set of copyrighted works from which any work in $D$ stems. We denote its complement $\mathcal{C}_{-D}=\mathcal{C}\setminus\mathcal{C}_{D}$.
@@ 7.2 Clean training algorithms: copy protection for blameless users in a clean room
7.2.1 The user’s clean-room distribution
@@
This is closely related to NAF. The model $p\leftarrow\mathsf{Train}(\mathsf{scrub}(D,c))$ is NAF’s safe model $\mathsf{safe}_{c}$ for an appropriately defined function $\mathsf{safe}$. The clean-room distribution $\tau_{-c}$ is the distribution over outputs that results when the user given $\mathsf{aux}$ interacts with $\mathsf{safe}_{c}$.
7.2.2 Clean-room blamelessness and copy protection
We postulate that blameless users can control their risk of copying when using a model trained in a clean room. Given $\beta>0$, a blameless user can guarantee that they copy a work $c$ that was “outside the clean room” with probability at most $\beta$. This should hold even if the user is exposed to $c$ in some other than through the model. People are constantly exposed to copyrighted works, yet we somehow to manage to avoid copying every day. Using a model trained without access to $c$ shouldn’t change that.
@@ 7.2.2 Clean-room blamelessness and copy protection
For $0\leq\beta\leq 1$, a user $\mathsf{u}$ is $\beta$-blameless in the clean room ($\beta$-blameless) with respect to $D$, $\mathcal{C}$, $\mathsf{Train}$ if for all $c\in\mathcal{C}$ and all $\mathsf{aux}$:55footnotemark: 5
@@
Otherwise, $\mathsf{u}$ is $\beta$-blameworthy. [^6] Equivalently, $\mathfrak{B}^{\mathsf{clean}}_{\beta}(D,\mathcal{C},\mathsf{ideas}):=\{\mathsf{u}:\forall c\in\mathcal{C},\forall\mathsf{aux},\tau_{-c}\bigl(\mathsf{SubSim}\bigl(\mathcal{C}_{-D}\cup\{c\}\bigr);~\mathsf{id}\bigr)\leq\beta\}.$
@@
Remark 7.7 (How small is $\beta$?).
We think of $\beta$ as the probability that a “truly blameless” user (whatever that means) nevertheless produces something substantially similar to a copyrighted work. There is a limit to how small this can be. Even a monkey on a typewriter—truly blameless if anyone is—has a non-zero $\beta$.
Getting a good estimate of achievable $\beta$ is beyond the scope of this work. It should be small enough to be very unlikely to occur by chance: $\beta\ll 1/100$ is certainly doable. It should be large enough to reflect that a relatively small amount of original expression is needed for copyright protection: $\beta\gg 1/10^{100}$ is impossible, even for the monkey. We conjecture that $\beta\approx 10^{-6}$ or $10^{-9}$ is achievable by a conscientious user.
7.2.3 Clean-room copy protection is not tainted
By Desideratum 2, a good definition of provable copy protection should exclude tainted training algorithms. The following theorem shows that clean-room copy protection passes this test, under a reasonable assumption on $\mathsf{ideas}$ and $\mathsf{SubSim}$. Roughly, the assumption is that there exist at least $n$ distinct ideas each of which can be expressed in $m\gg n$ ways satisfying a strong dissimilarity property: that for any distinct $w$ and $w^{\prime}$, the set of works substantially similar to both $w$ and $w^{\prime}$ is empty.
@@ 8 Differential privacy’s protection against copying
Many works have suggested that differential privacy [dwork2006calibrating] (DP) should protect against copying if the training data are deduplicated [bousquet2020synthetic, henderson2023foundation, VyasKB23, elkinkoren2024copyright, chen2024randomization, livni2024credit]. This section turns the suggestion into a theorem. For background on DP, see Appendix C.
At a high level, we show that DP training blameless users from copyright infringement (in the sense of Definition 7.6) if the training dataset is golden, a copyright-deduplication condition defined below. If so, the probability of copying is at most $\approx e^{\varepsilon}\beta N_{D}$, where $\beta$ is a user’s probability of copying in the clean room and $N_{D}=|\mathcal{C}_{D}|$ is the number of copyrighted works to which $D$ grants access. To guarantee the risk is at most $\kappa$, the user sets $\beta\approx{\kappa}/{e^{\varepsilon}N_{D}}$. The linear dependence on $N_{D}$ is not ideal, but may sometimes be good enough or improved (Remark 8.4.)
This allows a user choose their appetite for copyright risk $\kappa$ and then act to set $\beta$ accordingly. Stephen King writing his next bestseller has a much lower risk tolerance ($\kappa=10^{-9}$) than Joe Schmoe writing a wedding toast ($\kappa=10^{-1}$). Most users are somewhere in between. For $\varepsilon=5$ and $N_{D}=200,000$ Stephen King needs $\beta\approx 10^{-15}$ while Joe Schmoe only needs $\beta\approx 10^{-6}$. If $\beta\approx 10^{-15}$ is too small (Remark 7.7), Stephen King can instead forgo using the LLM altogether.
A dataset $D$ is golden if at most one item $w\in D$ stems from any copyrighted work cc. [^7] We adapt and formalize this idea from [VyasKB23], where it appears as an informal condition that suffices for their analysis: “It is important that when we omit a data point $x$ it does not share copyrighted content with many other data points that were included in the training set.” We also borrow the term “golden dataset”, though [VyasKB23] uses it to refer to something entirely different: A dataset that is “carefully scrutinized to ensure that all material in it is not copyrighted or [is] properly licensed.” That condition doesn’t suffice, as even licensed derivates could cause unlicensed copying by a generative model. The same is true for NAF, despite their suggestion that it would yield a safe model.   That item could be $c$ itself or a derivative work. No protected original expression appears in more than one element of a golden dataset. For example, it can contain one copy or parody of the Abbey Road album cover (not both), and many parodies of the out-of-copyright Mona Lisa. Remark 8.5 explains why golden datasets will typically remain golden as the set of in-copyright works evolves over time.
@@
Let $G=(\mathcal{W},E)$ be a copyright dependency graph, and $\mathcal{C}\subseteq\mathcal{W}$ be the set of in-copyright works. A dataset $D$ is golden (with respect to $G$, $\mathcal{C}$) if for all $c\in\mathcal{C}$ there is at most one $w\in D$ such that $(c,w)\in E$.
@@
Fix $\mathcal{C}\subseteq\mathcal{W}$, $D$, and user $\mathsf{u}$ that is $\beta$-blameless (Def. 7.5). We must show that for all $\mathsf{aux}$, $\tau(\mathsf{SubSim}(\mathcal{C});~\mathsf{aux})\leq\kappa$.
Because $D$ is golden, $|D\setminus\mathsf{scrub}(D,c)|\leq 1$. Hence, datasets $\mathsf{scrub}(D,c)$ and $D$ are either equal or neighboring for all $c\in\mathcal{C}$. Applying DP and post-processing, we have $\tau(E)\leq e^{\varepsilon}\cdot\tau_{-c}(E)+\delta$ for all events $E$ and all $c\in\mathcal{C}$.
$\tau(\mathsf{SubSim}(\mathcal{C});~\mathsf{aux})$ $\leq\tau(\mathsf{SubSim}(\mathcal{C}_{-D});~\mathsf{aux})+\sum_{c\in\mathcal{C}_{D}}\tau(\mathsf{SubSim}(c);~\mathsf{aux})$ (union bound)
$\leq\beta+N_{D}\delta+e^{\varepsilon}\cdot\sum_{c\in\mathcal{C}_{D}}\tau_{-c}(\mathsf{SubSim}(c);~\mathsf{aux})$ (DP; Prop. 8.3 below)
$\leq(e^{\varepsilon}N_{D}+1)\beta+N_{D}\delta$ ($\beta$-blameless)
In the proof above, we need to bound the probability $\tau(\mathsf{SubSim}(\mathcal{C}_{-D});~\mathsf{aux})$ of similarity with some work in $\mathcal{C}_{-D}$ in the real distribution $\tau$. Blamelessness only gives us a bound in the clean-room distribution $\tau_{-c}$. Proposition 8.3 shows that the latter implies the former.
@@
Let $\mathsf{u}$ be $\beta$-blameless w.r.t. $D$, $\mathcal{C}$, $\mathsf{Train}$. For all $\mathsf{aux}$: $\tau\left(\mathcal{C}_{-D};~\mathsf{aux}\right)\leq\beta.$
@@
Remark 8.4 (On the linear dependence on $N_{D}$).
Notice that $\kappa>N_{D}\cdot(\beta+\delta)$ grows linearly with $N_{D}=|\mathcal{C}_{D}|$. This only yields a meaningful bound if $N_{D}\ll 1/(\beta+\delta)$. Along with the difficulty of creating enormous golden datasets (Section 9), this is another reason that our approach is most practical when only medium-sized datasets are required.
On the other hand, the factor of $N_{D}=|\mathcal{C}_{D}|$ from the union bound is unreasonably pessimistic. Ideas sometimes differ so greatly that the risk of substantial similarity is essentially 0: a photorealistic bird will never resemble Dr. Seuss’s Cat in the Hat. If the user is targeting a fixed set of ideas $\mathsf{id}$ (conditioning $\tau$ on $\mathsf{id}$), one might instead take the union bound over $\mathcal{C}_{D}^{\mathsf{id}}=\{c\in\mathcal{C}_{D}:\mathsf{id}\in\mathsf{ideas}(\mathsf{SubSim}(c))\}$. We expect that typically $N_{D}^{\mathsf{id}}:=|\mathcal{C}_{D}^{\mathsf{id}}|\ll|\mathcal{C}_{D}|=N_{D}$.
@@
9 Discussion
9.1 Indemnification for genAI outputs